# Quantum Blackwell’s Ordering and Differential Privacy

### v1

@@ V-A Main characterization
[…]
@@
[…]
Motivated by the close relation between privacy and hypothesis testing in the lemma below, we give an equivalent condition for a pair of quantum states to be $(\varepsilon,\delta)$-DP.
@@
A pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ (see Definition 16) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1)$ if and only if the following holds,
$D^{\alpha}_{H}{(\rho\|\sigma)}\leq-\log f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1].$ (56)
where $f_{\varepsilon,\delta}(\alpha):=\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\alpha),0\right\}$ for any $\alpha\in[0,1]$.
@@
Note that the above version of quantum $(\varepsilon,\delta)$-DP implies that if for each $\alpha\in[0,1]$ $D^{\alpha}_{H}(\rho^{\prime}\|\sigma^{\prime})=-\log f_{\varepsilon,\delta}(\alpha)$ for some pair of quantum states $\rho^{\prime},\sigma^{\prime}$ over any arbitrary finite dimension Hilbert space, then for any pair of quantum states $\rho,\sigma\in\mathfrak{D}_{(\varepsilon,\delta)}$ (see Definition 16) is at least as hard as distinguishing between the pair $(\rho^{\prime},\sigma^{\prime})$. The above intuition provides a notion of the weakest (most informative) quantum $(\varepsilon,\delta)$-DP pairs of quantum states, which are the least “hard to differentiate”. We formalize this notion in the definition below.
@@
$D^{\alpha}_{H}{(\rho\|\sigma)}=-\log f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1].$ (57)
where $f_{\varepsilon,\delta}(\alpha):=\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\alpha),0\right\}$ for any $\alpha\in[0,1]$.
[…]
Theorem 2.
[…]
where $(\rho^{\prime},\sigma^{\prime})$ is some pair of quantum states which is the weakest (most informative) $(\varepsilon,\delta)$-DP (see Definition 19).
Proof.
If a pair $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, then, from Lemma 3 and the fact that $D^{\alpha}_{H}(\rho^{\prime}\|\sigma^{\prime})=-\log f_{\varepsilon,\delta}$ for every $\alpha\in[0,1]$, the following holds,
$D^{\alpha}_{H}(\rho\|\sigma)\leq D^{\alpha}_{H}(\rho^{\prime}\|\sigma^{\prime}),\forall\alpha\in[0,1].$ (59)
[…]
In the lemma below, we show the existence of a pair of quantum states that is the weakest (most informative) quantum $(\varepsilon,\delta)$-DP (see Definition 19) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$ in the Blackwell sense.
@@
Consider the quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ mentioned in (51) and (52), for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$. Then, we have $D^{\alpha}_{H}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})=-\log\max\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\delta-\alpha),0\}$ for any $\alpha\in[0,1]$.
@@
V-B Blackwell Dominance under Quantum Differential Privacy
From Lemmas 3 and 4 and Definition 15, we note that any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, gets dominated by the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$. Formally, it implies the following.
@@ V-B Blackwell Dominance under Quantum Differential Privacy
$\left\{(\rho,\sigma)\right\}\preceq_{B_{Q}}\left\{(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})\right\}.$ (60)
@@
  • Note that the quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ mentioned in Lemma 4 satisfies weakest (most informative) $(\varepsilon,\delta)$-DP condition for fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$ and therefore, they have a fixed-point at $\alpha=\frac{1-\delta}{e^{\varepsilon}+1}$.
  • For $\delta=0$ and any $\varepsilon\geq 0$, we define two quantum states $\rho_{\varepsilon}:=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{1}{1+e^{\varepsilon}}|1\rangle\langle 1|$ and $\sigma_{\varepsilon}:=\frac{1}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|1\rangle\langle 1|$. Observe that, it can be verified that $\rho_{\varepsilon}$ and $\sigma_{\varepsilon}$ are weakest (most informative) $(\varepsilon,0)$-DP states.
@@
Corollary 2.
[…]
The reference [19, Theorem 18] showed Corollary 2 in the classical case.
Corollary 3.
[…]
where $\rho_{(\varepsilon)}:=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{1}{1+e^{\varepsilon}}|1\rangle\langle 1|$ and $\sigma_{(\varepsilon)}:=\frac{1}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|1\rangle\langle 1|$.
Now using Corollary 2, for any pair of quantum states $\rho,\sigma\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, we can derive upper-bounds in terms of quantum hockey-stick divergence. However, it is important to note that for $\delta>0$, $\operatorname{supp}(\rho_{(\varepsilon,\delta)})\not\subseteq\operatorname{supp}(\sigma_{(\varepsilon,\delta)})$. Therefore, the divergence that requires support inclusion is not well-defined for the pair of quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$. Although for $\delta=0$, using Corollary 3, since $\operatorname{supp}(\rho_{\varepsilon})=\operatorname{supp}(\sigma_{\varepsilon})$, for any $\varepsilon,0$-DP quantum state pairs, we can derive upper-bounds on quantum relative entropy, quantum Rényi divergence in the corollary below.
@@
    $D_{\alpha}(\rho\|\sigma)\leq\frac{1}{\alpha-1}\left[\log\left(e^{\alpha}\varepsilon+e^{(1-\alpha)\varepsilon}\right)-\log\left(1+e^{\varepsilon}\right)\right]$,  for all $\alpha\in[0,1]\cup(1,2)$ ,
  3. (iii)
    ${\left\lVert\rho-\sigma\right\rVert}_{1}\leq\varepsilon\tanh\left(\frac{\varepsilon}{2}\right)$.
@@
(i) From Fact 11 it follows that $D(\cdot\|\cdot)$ is a valid divergence and thus, we can write the following,
$D(\rho^{\prime}\|\sigma^{\prime})$ $=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\varepsilon-\frac{1}{1+e^{\varepsilon}}\varepsilon$
$=\varepsilon\left(\frac{e^{\frac{\varepsilon}{2}}-e^{-\frac{\varepsilon}{2}}}{e^{\frac{\varepsilon}{2}}+e^{-\frac{\varepsilon}{2}}}\right)$
$=\varepsilon\tanh\left(\frac{\varepsilon}{2}\right).$ (63)
Therefore, (61) of Corollary 2 completes the proof of 1) of Corollary 4.
(ii) From Fact 13 it follows that for $\alpha\in[0,1]\cup(1,2]$, $D_{\alpha}(\cdot\|\cdot)$ is a valid divergence and thus, we can write the following,
$D_{\alpha}(\rho^{\prime}\|\sigma^{\prime})$ $=\frac{1}{\alpha-1}\log\left(\left(\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\right)^{\alpha}\left(\frac{1}{1+e^{\varepsilon}}\right)^{1-\alpha}+\left(\frac{1}{1+e^{\varepsilon}}\right)^{\alpha}\left(\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\right)^{1-\alpha}\right)$
$=\frac{1}{\alpha-1}\log\left(\frac{e^{\alpha}\varepsilon+e^{(1-\alpha)\varepsilon}}{1+e^{\varepsilon}}\right)$
$=\frac{1}{\alpha-1}\left[\log\left(e^{\alpha}\varepsilon+e^{(1-\alpha)\varepsilon}\right)-\log\left(1+e^{\varepsilon}\right)\right].$ (64)
Therefore, (61) of Corollary 2 completes the proof of 2) of Corollary 4.
(iii) As ${\left\lVert\mathcal{N}(\rho)-\mathcal{N}(\sigma)\right\rVert}_{1}=2E_{1}(\mathcal{N}(\rho)\|\mathcal{N}(\sigma))$, from Fact 14 it follows that ${\left\lVert\cdot-\cdot\right\rVert}_{1}$ is a valid divergence and thus, we can write the following,
${\left\lVert\rho^{\prime}-\sigma^{\prime}\right\rVert}_{1}$ $=\left|\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}\right|+\left|\frac{1-e^{\varepsilon}}{1+e^{\varepsilon}}\right|$
$\overset{a}{=}2\left|\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}\right|$
$=2\tanh\left(\frac{\varepsilon}{2}\right).$ (65)
Therefore, (61) of Corollary 2 completes the proof of 3) of Corollary 4.
VI Quantum Privatized Parameter Inference
In the context of randomized response, a fundamental challenge is balancing the privacy of respondents with the statistical utility of the information gathered by an investigator. Let’s consider a scenario where binary information, represented by $\{0,1\}$, is encoded into two quantum states, $\rho$ and $\sigma$. The privacy of this encoding is quantified by the $(\varepsilon,\delta)$-DP constraint on the pair $(\rho,\sigma)$.
We assume that the behavior of the respondents is considered as random sampling with replacement from the investigator’s perspective, and the investigator is interested only in the ratio between 0 and $1$ so that the underlying statistical information $\theta$ follows a binomial distribution $p_{\theta}$, where $p_{\theta}(0)=\theta$ and $p_{\theta}(1)=1-\theta$. That is, the key task is the inference of the parameter $\theta$ by the investigator.
With the quantum encoding, this translates to a parametrized state $\rho_{\theta}:=\theta\rho+(1-\theta)\sigma$, where $\theta\in[0,1]$. If the investigator collects $n$ responses, the composite state is $\rho_{\theta}^{\otimes n}$.

### v3

@@ IV-A Quantum ordering of informativeness
Then, for any $f\in\mathcal{F}$ (see Definition 10) we have,
$D_{f}(\rho_{1}\|\rho_{2})\geq D_{f}(\sigma_{1}\|\sigma_{2}),$
whenever $D_{f}$ is well-defined (see Definition 10) for the pairs $(\rho_{1},\rho_{2})$ and $(\sigma_{1},\sigma_{2})$.
Before proceeding with the proof of Theorem 1, we first state the following lemma, which will be required in proving Theorem 1.
Lemma 1.
then,
$E_{\gamma}(\rho_{1}\|\rho_{2})$ $\geq E_{\gamma}(\sigma_{1}\|\sigma_{2}),\forall\gamma\geq 0,$ (54)
$D_{\tilde{\alpha}}(\rho_{1}\|\rho_{2})$ $\geq D_{\tilde{\alpha}}(\sigma_{1}\|\sigma_{2}),\forall\tilde{\alpha}\in[0,1)\cup(1,+\infty),$ (55)
$D(\rho_{1}\|\rho_{2})$ $\geq D(\sigma_{1}\|\sigma_{2}),$ (56)
@@
See Appendix -A for the proof.
Using the above lemma, we now prove Theorem 1 using the following series of inequalities,
$D_{f}(\rho_{1}\|\rho_{2})$ $\overset{(a)}{=}\int_{0}^{\infty}f^{\prime\prime}(\gamma)E_{\gamma}(\rho_{1}\|\rho_{2})d\gamma$
$\overset{(b)}{\geq}\int_{0}^{\infty}f^{\prime\prime}(\gamma)E_{\gamma}(\sigma_{1}\|\sigma_{2})d\gamma$
$\overset{(c)}{=}D_{f}(\sigma_{1}\|\sigma_{2}),$
where $(a)$ and $(c)$ follows from Fact 5 and $(b)$ follows from (54) of Lemma 1 and the fact that the double derivative of $f$ exists and $f$ is a convex function. This completes the proof of Theorem 1.
Remark 3.
We note here that (54) of Lemma 1 was also proven in [5, Theorem 22].
Remark 4.
There have been various notions of ordering of informativeness in the quantum case in the earlier works, see for example, [28, 8, 21, 24, 9, 10, 7, 33]. However, it is not very clear whether these orderings are useful for characterizing differentially private mechanisms, which is the main aim of this manuscript.
IV-B Differential privacy: A quantum hypothesis testing perspective
In this subsection, we consider differential privacy from a quantum hypothesis testing perspective. The main goal of quantum differential privacy is to ensure that a quantum mechanism produces output states that are nearly indistinguishable when applied to neighboring quantum inputs. Formally, for two neighboring quantum inputs producing output states $\rho,\sigma$, the distinguishability between $\rho$ and $\sigma$ must be limited by the privacy parameters. Intuitively, this means that even if an investigator (adversary) knows the mechanism and observes its output, the investigator should not be able to reliably infer which quantum input (respondent) generated it. More formally, with high probability, no hypothesis test—regardless of the investigator’s strategy—can reliably infer the respondent’s individual contribution from the output.
Motivated by the close relation between privacy and hypothesis testing, in the lemma below, we give an equivalent condition for a pair of quantum states to be $(\varepsilon,\delta)$-DP (see Definition 1).
@@ IV-B Differential privacy: A quantum hypothesis testing perspective
Consider any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ (see Definition 1) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1)$. Then, the following holds,
$D^{\alpha}_{H}{(\rho\|\sigma)}\leq-\log f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1],$ (57)
or, equivalently,
$\beta(\alpha|\rho\|\sigma)\geq f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1],$ (58)
where,
$f_{\varepsilon,\delta}(\alpha):=\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\delta-\alpha),0\right\},$ (59)
for any $\alpha\in[0,1]$.
@@
Note that the above version of quantum $(\varepsilon,\delta)$-DP implies that if for each $\alpha\in[0,1]$ $D^{\alpha}_{H}(\rho^{\prime}\|\sigma^{\prime})=-\log f_{\varepsilon,\delta}(\alpha)$ for some pair of quantum states $\rho^{\prime},\sigma^{\prime}$ over any arbitrary finite dimensional Hilbert space, then for any pair of quantum states $\rho,\sigma\in\mathfrak{D}_{(\varepsilon,\delta)}$ (see Definition 1) is at least as hard as distinguishing between the pair $(\rho^{\prime},\sigma^{\prime})$. The above intuition provides a notion of the weakest (most informative) quantum $(\varepsilon,\delta)$-DP pairs of quantum states, which are the least “hard to differentiate”. We formalize this notion in the definition below.
@@
$D^{\alpha}_{H}{(\rho\|\sigma)}=-\log f_{\varepsilon,\delta}(\alpha),\forall\alpha\in[0,1],$ (60)
In the lemma below, we show the existence of a pair of quantum states that is the weakest (most informative) quantum $(\varepsilon,\delta)$-DP (see Definition 18) for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$.
@@
Consider the quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ mentioned in (47) and (48), for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$. Then, we have $D^{\alpha}_{H}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})=-\log f_{\varepsilon,\delta}(\alpha)$ for any $\alpha\in[0,1]$.
@@
Therefore, for a pair $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ with some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, Lemmas 2 and 3 guarantee the relation,
$D^{\alpha}_{H}(\rho\|\sigma)$ $\leq-\log f_{\varepsilon,\delta}(\alpha)=D^{\alpha}_{H}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (61)
for every $\alpha\in[0,1]$.
Further, from Lemmas 2 and 3 and Definition 17, we note that any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, gets dominated by the pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$. Formally, it implies the following.
@@
$\left\{\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)}\right\}\succeq_{Q}\left\{\rho,\sigma\right\}.$ (62)
@@
  • Note that the quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ mentioned in Lemma 3, which satisfies weakest (most informative) $(\varepsilon,\delta)$-DP condition for fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, has a fixed-point at $\alpha=\frac{1-\delta}{e^{\varepsilon}+1}$.
  • For $\delta=0$ and any $\varepsilon\geq 0$, we define two quantum states,
    $\rho_{(\varepsilon)}$ $:=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{1}{1+e^{\varepsilon}}|1\rangle\langle 1|,$ (63)
    $\sigma_{(\varepsilon)}$ $:=\frac{1}{1+e^{\varepsilon}}|0\rangle\langle 0|+\frac{e^{\varepsilon}}{1+e^{\varepsilon}}|1\rangle\langle 1|,$ (64)
    Observe that it can be verified that $\rho_{(\varepsilon)}$ and $\sigma_{(\varepsilon)}$ are the weakest (most informative) $\varepsilon$-DP pair of quantum states.
We now state the main theorem of this section, which provides a bound on $D_{f}(\rho\|\sigma)$ (see Definition 10) for any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$.
Theorem 2.
If a pair of quantum states $(\rho,\sigma)$ belongs to $\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, then for any $f\in\mathcal{F}$, we have,
$D_{f}(\rho\|\sigma)\leq D_{f}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (65)
whenever $D_{f}$ is well-defined for the pairs $(\rho_{1},\rho_{2})$ and $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$. Also, for any $\gamma\geq 0$ and $\alpha\in[0,1)\cup(1,+\infty)$, we have,
$E_{\gamma}(\rho\|\sigma)$ $\leq E_{\gamma}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (66)
$D_{\alpha}(\rho\|\sigma)$ $\leq D_{\alpha}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (67)
$D(\rho\|\sigma)$ $\leq D(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (68)
$\check{D}_{\alpha}(\rho\|\sigma)$ $\leq\check{D}_{\alpha}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}).$ (69)
Proof.
Suppose a pair $(\rho,\sigma)$ belongs to $\mathfrak{D}_{(\varepsilon,\delta)}$ for some fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$. Then, the combination of (61) and Theorem 1 implies the relation $D_{f}(\rho\|\sigma)\leq D_{f}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})$. The combination of (61) and Lemma 1 also implies (66), (67), and (68). Finally, using Fact 7, we have,
$\check{D}_{\alpha}(\rho\|\sigma)\leq D_{\alpha}(\rho\|\sigma)\leq D_{\alpha}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)})=\check{D}_{\alpha}(\rho_{(\varepsilon,\delta)}\|\sigma_{(\varepsilon,\delta)}),$ (70)
where the final equality follows from Fact 8, since $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$ are commutative with each other. This proves Theorem 2.
Further, for $\delta=0$, the theorem below gives an upper-bound on $D_{f}$ (for any $f\in\mathcal{F}$) between any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,0)}$.
Corollary 2.
If any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,0)}$ for some fixed $\varepsilon\geq 0$, then for any $f\in\mathcal{F}$, $\gamma\geq 0$ and $\alpha\in[0,1)\cup(1,+\infty)$, we have,
$D_{f}(\rho\|\sigma)$ $\leq D_{f}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)}),$ (71)
$E_{\gamma}(\rho\|\sigma)$ $\leq E_{\gamma}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)}),$ (72)
$D_{\alpha}(\rho\|\sigma)$ $\leq D_{\alpha}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)}),$ (73)
$\check{D}_{\alpha}(\rho\|\sigma)$ $\leq\check{D}_{\alpha}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)}).$ (74)
whenever $D_{f}$ is well-defined for the pairs $(\rho,\sigma)$ and $(\rho_{(\varepsilon)},\sigma_{(\varepsilon)})$ (defined in (63) and (64) respectively).
Remark 6.
Note that for some fixed $\varepsilon\geq 0$, the pair $(\rho_{(\varepsilon)},\sigma_{(\varepsilon)})$ is commutative and $D_{\max}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)})=D_{\max}(\sigma_{(\varepsilon)}\|\rho_{(\varepsilon)})$ $\varepsilon$. Further, Definition 1 directly guarantees that any pair of quantum states $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,0)}$ satisfies
$D_{\max}(\rho\|\sigma)$ $\leq\varepsilon=D_{\max}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)}),$
$D_{\max}(\sigma\|\rho)$ $\leq\varepsilon=D_{\max}(\sigma_{(\varepsilon)}\|\rho_{(\varepsilon)}).$
Therefore, from Fact 11, there exists a CP-TP map $\mathcal{E}$, such that $\rho=\mathcal{E}(\rho_{(\varepsilon)})$ and $\sigma=\mathcal{E}(\sigma_{(\varepsilon)})$. Hence, Definition 14 implies
@@
It is important to note that for $\delta>0$, $\operatorname{supp}(\rho_{(\varepsilon,\delta)})\not\subseteq\operatorname{supp}(\sigma_{(\varepsilon,\delta)})$. Therefore, the divergence that requires support inclusion is not well-defined for the pair of quantum states $\rho_{(\varepsilon,\delta)}$ and $\sigma_{(\varepsilon,\delta)}$. Although for the case when $\delta=0$, since $\operatorname{supp}(\rho_{\varepsilon})=\operatorname{supp}(\sigma_{\varepsilon})$, for any $\varepsilon$-DP quantum state pairs, using Corollary 2, we derive upper-bounds on quantum relative entropy and $L_{1}$ distance as mentioned in the corollary below.
@@
    ${\left\lVert\rho-\sigma\right\rVert}_{1}\leq 2\tanh\left(\frac{\varepsilon}{2}\right)$.
@@
(i) Fact 6 yields that $D(\cdot\|\cdot)=D_{f}(\cdot\|\cdot)$, where $f(x)=x\log x$. Further, the pair $(\rho_{(\varepsilon)},\sigma_{(\varepsilon)})$ satisfies the following,
$D(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)})$ $=\frac{e^{\varepsilon}}{1+e^{\varepsilon}}\varepsilon-\frac{1}{1+e^{\varepsilon}}\varepsilon$
$=\varepsilon\left(\frac{e^{\frac{\varepsilon}{2}}-e^{-\frac{\varepsilon}{2}}}{e^{\frac{\varepsilon}{2}}+e^{-\frac{\varepsilon}{2}}}\right)=\varepsilon\tanh\left(\frac{\varepsilon}{2}\right).$ (75)
Therefore, (71) of Corollary 2 completes the proof of (i) of Corollary 3.
(ii) The pair $(\rho_{(\varepsilon)},\sigma_{(\varepsilon)})$ satisfies the following,
${\left\lVert\rho_{(\varepsilon)}-\sigma_{(\varepsilon)}\right\rVert}_{1}$ $=\left|\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}\right|+\left|\frac{1-e^{\varepsilon}}{1+e^{\varepsilon}}\right|$
$\overset{(a)}{=}2\left|\frac{e^{\varepsilon}-1}{1+e^{\varepsilon}}\right|=2\tanh\left(\frac{\varepsilon}{2}\right).$ (76)
where $(a)$ follows from the fact that $e^{\varepsilon}\geq 1$ as $\varepsilon\geq 0$. Note that we have ${\left\lVert\rho-\sigma\right\rVert}_{1}=2\left|\rho-\sigma\right|_{+}=2E_{1}(\rho\|\sigma)$, where the last equality follows from Fact 10. Therefore, the relation (72) of Corollary 2 derives that ${\left\lVert\rho-\sigma\right\rVert}_{1}=2E_{1}(\rho\|\sigma)\leq 2E_{1}(\rho_{(\varepsilon)}\|\sigma_{(\varepsilon)})={\left\lVert\rho_{(\varepsilon)}-\sigma_{(\varepsilon)}\right\rVert}_{1}$. This completes the proof of (ii) of Corollary 3.
V Quantum Privatized Parameter Inference
V-A Problem setting
We consider a statistical inference problem with $n$ respondents with the privacy setting. Respondent $i\in\{1,\ldots,n\}$ holds a private bit $X_{i}\in\{0,1\}$, and the investigator aims to infer a population parameter $\theta\in[0,1]$ that governs the relative frequency of the two outcomes. Concretely, since (from the investigator’s perspective) the respondents’ behavior is modeled as random sampling with replacement, the $n$ private bits can be treated as i.i.d. draws from a Bernoulli law $p_{\theta}$ on $\{0,1\}$ with $p_{\theta}(0)=\theta$ and $p_{\theta}(1)=1-\theta$. Equivalently, the investigator treats the collected answers as random sampling with replacement from a population with composition parameterized by $\theta$, so that $(X_{1},\ldots,X_{n})$ is (approximately) i.i.d. under $p_{\theta}$. The key task is to estimate $\theta$ from the $n$ (privatized) responses.
To gather information while protecting privacy, each respondent releases not the raw bit $X_{i}$ but a quantum system whose state depends on $X_{i}$. Specifically, we fix two quantum states $\rho$ and $\sigma$, and respondent $i$ releases state $\rho$ if $X_{i}=0$ and state $\sigma$ if $X_{i}=1$. Thus, each response is produced by an encoding map $x\in\{0,1\}\mapsto\rho_{x}$, where $\rho_{0}:=\rho$ and $\rho_{1}:=\sigma$. The privacy requirement is imposed at the level of each respondent’s released system. Namely, we require that the pair of possible outputs of an individual respondent, $(\rho,\sigma)$, satisfies $(\varepsilon,\delta)$-DP in the sense of Definition 1, i.e., $(\rho,\sigma)\in\mathfrak{D}_{(\varepsilon,\delta)}$.
Parametrized states induced by the population model. Under the Bernoulli model with parameter $\theta$, a single privatized response is described by the mixture $\rho_{\theta}:=\theta\,\rho+(1-\theta)\,\sigma$ with $\theta\in[0,1]$. Since the investigator collects $n$ responses, the resulting joint state is modeled as the product state $\rho_{\theta}^{\otimes n}$. This is the quantum analogue of observing $n$ i.i.d. privatized samples drawn according to $p_{\theta}$, where the unknown parameter $\theta$ controls the mixture proportions between $\rho$ and $\sigma$.
We also define $\rho_{(\varepsilon,\delta),\theta}$ analogously when the respondents encode their information using the weakest/most informative $(\varepsilon,\delta)$-DP pair $(\rho_{(\varepsilon,\delta)},\sigma_{(\varepsilon,\delta)})$ constructed earlier in the paper. That is, $\rho_{(\varepsilon,\delta),\theta}:=\theta\,\rho_{(\varepsilon,\delta)}+(1-\theta)\,\sigma_{(\varepsilon,\delta)}$. This reference pair plays a canonical role in our analysis: since it is the most informative among all $(\varepsilon,\delta)$-DP pairs, it provides the fundamental benchmark for the maximum inferential power (e.g., hypothesis testing performance and Fisher information) achievable by any privatized mechanism under the same privacy parameters.